plotted on the Celsius scale, the straight ine can be volume is zero and the temperature is-273°C the ideal gas law can be Part C so that the gas pressure changed from 2.00 L to 3.90 L 7 (The mass of helium added 3.90 L V 1.00
Answer #1

molar mass of He = 4 g/mol

initial number of moles of He, ni = mass / molar mass

= 2.00/4

= 0.50 mol

Now calculate the final number of moles of He

Given:

ni = 0.50 mol

Vi = 2.00 L

Vf = 3.90 L

use:

Vi/ni = Vf/nf

2.0 L / 0.50 mol = 3.9 L / nf

nf = 0.975 mol

final mass of He = nf*molar mass of He

= 0.975*4

= 3.9 g

mass of He added = final mass - initial mass

= 3.9 g - 2.00 g

= 1.9 g

Answer: 1.9 g